OpenStudy (anonymous):

-3b+8=2b+48 Plz help I don't think I am doing this right

OpenStudy (anonymous):

-3b+8=2b+48 -2 -2 you want to get b by itself so to do that you need to subtract 2 from both sides -5b+8=48 -8 -8 -5b=40 now to get b by itself you divide both sides by

OpenStudy (anonymous):

b=-8

OpenStudy (kewlgeek555):

Well, to see if this is true... 8 - (-3) = b 48 - 2 = b If those two bs are the same, then you have the right answer, if not, I think ti would be called a "no solution".

OpenStudy (anonymous):

Thank u guys so much u really helped :)

OpenStudy (anonymous):

No problem :)